OpenStudy (anonymous):

a4(a-1)(a3) (#'s are exponents).. A. a7 B. a6 C. a-12

OpenStudy (anonymous):

he exponents. So #1 is 3^-3 m^-4 t^11 or like you said, t^11 / (3^3 m^4) #2: 2.7 ÷ 9 = 0.3 so 0.3 (10 ^ -4) = 0.00003 # just becomes c / (a^4b^2) 4 is 4^3 m^6 n^-12 or 64m^6 / n^12 Then on the other, continue by grouping: a^3 + a^2 = a^2(a + 1) -a - 1 = -1(a + 1) so that part becomes (a + 1)(a^2 - 1) which then becomes (a + 1)(a + 1)(a - 1) so the final factored form is 7(a + 1)(a + 1)(a - 1) 2 months ago

OpenStudy (anonymous):

a7 is wrie